OpenStudy (anonymous):

How do you make a line of best fit for a sideways parabola? This is part of a Physics assignment (our teacher wants us to graph using special physics graph rules, so we are doing lots of graphing)

OpenStudy (anonymous):

And she was telling us about how we need to straighten the parabola to find the line of best fit, using some variable relationship

OpenStudy (aum):

One simple example of a sideways parabola is: x = ay^2 Solving for y we get: \(\large y = \sqrt{\frac xa} = \frac{1}{\sqrt{a}} * \sqrt{x}\). Instead of plotting y against x, you can plot y against \(\large \sqrt{x}\) and you will get a straight line.

OpenStudy (anonymous):

The formula my teacher gave us was y^2=mx+b

OpenStudy (anonymous):

And I believe that she wanted us to use this. She also wanted us to find it using some sort of variable relationship, like y^2 vs x